solve for x
7/(4x) = 5/(6x)-3

Found 2 solutions by dabanfield, rfadrogane:
Answer by dabanfield(803)

solve for x
7/(4x) = 5/(6x)-3
To get rid of the denominators we need to multiply all terms by a common multiple of 4x and 6x. 12x is the smallest such multiple.
So we multiply both sides of the equation by 12x:
(12x)*7/(4x) = (12x)*5/(6x) - (12x)*3 =
(12x/4x)*7 = (12x/6x)*5 - 36x
3*7 = 2*5 - 36x
36x = 10-21
36x = -11
x=-11/36

Answer by rfadrogane(214)

solve for x
7/(4x) = 5/(6x)-3
Sol'n:
7[6x-3] = 5[4x]
42x-21 = 20x
12x = 21
x = 21/12
x = 7/4----answer